# Confidence Driven Classification of Application Types in the Presence of Background Network Traffic

### v3

@@ I Introduction
It is undesirable for background traffic to be labeled as an application traffic by a classifier, or for relevant application traffic to be incorrectly identified as background traffic. These misclassifications reduce both the accuracy and coverage needed for monitoring the usage, performance, and resource needs of different applications. In such scenarios, instead, it may be preferable that the deep learning model assign labels to traffic flow only when it can do so with high confidence in accuracy. Therefore, in this paper, we develop a framework that can characterize the confidence of a model’s classification decision in a manner that significantly improves the classification performance. Our main innovations in this paper are as follows:
@@
    We investigate the trade-off between accuracy and coverage in traffic classification and subsequently design a classification framework with well-calibrated confidence scores that can help achieve different operating points in this trade-off based on the confidence threshold.
@@ III Deep Learning Classifiers: State of the Art
Various feature representations have been proposed for deep learning models. One of the earliest approaches simply fed the first $n$ raw bytes of each flow directly into a model for malware detection [21]. Building on this idea, another study [25] converted the first 512 bytes (beginning at the transport layer header) across 32 packets per flow into an image‐based representation for classification. As an alternative to feeding raw bytes of a flow directly into the model, ET-BERT [6] segmented traffic into “BURST,” which is a sequence of consecutive packets traveling in the same direction. Each BURST’s raw datagram bytes were split into two-byte bigrams, and the bigrams were converted to tokens for model input [6]. The model learned the contextual relationships among these tokens through self-supervised pre-training and was subsequently fine-tuned for different classification tasks [6]. However, one prior work has found that classifiers trained on raw packet payload bytes in encrypted traffic may overfit on unencrypted handshake fields. In their evaluation, masking SNI alone caused accuracy of the classifier to fall to 38.19%, and masking all handshake fields drove it down to 27.52% [10]. Supporting this concern, a different study [22] demonstrated that ET-BERT may overfit on specific packet-level information, ranging from SNI to session-specific header fields, when the dataset is not carefully curated. These findings highlight the importance of selecting robust features to ensure the model does not depend on specific byte patterns that vary across traffic.
In contrast, other studies employed time series features that are not dependent on specific packet bytes, such as sequences of packet sizes, arrival times, and directions for traffic classification [16, 7, 1]. Since these features are universally available [1] and help prevent overfitting on packet level bytes, we experimentally evaluate classifiers that use time series features derived from sequences of $n$ packets in TCP and UDP traffic. [^2] In this paper, we group UDP packets sharing the same 5-tuple and refer to them collectively as a UDP flow. We use Flow Sequence Network (FS-Net) [7] and our implementation of a Bidirectional LSTM (BiLSTM) [13], as FS-Net and various forms of LSTM have been shown to perform effectively on time series features in prior studies [1, 11, 23, 7, 6]. Each model is trained with cross-entropy loss, using a distinct combination of time series features and data balancing techniques. Although FS-Net overall achieves better performance, our BiLSTM yields comparable results, demonstrating that performance trends persist across different architectures, feature sets, and data balancing methods. Hence, we include both models in our evaluation. The architecture of our BiLSTM is as follows:
@@
We use the time series features defined in [1] to train the BiLSTM. Each flow is represented by a 40 × 2 feature vector corresponding to the first 40 packets, where at each index, the two features are: packet arrival time relative to the flow start and packet size. To encode directionality, we negate the packet size of the server packets. Since we have more web browsing flows compared to other application types, we use data augmentation to balance the classes to have the same amount of flows in the training set, since appropriate data augmentation improves training [20]. We adapt the translation technique that works well with time series features [20] and derive an augmented feature vector from a sampled original one as follows: we randomly select a start index of a subsequence and shift it either to the left or to the right by n steps. For left shifts, if the original traffic flow had more than 40 packets, we use the subsequent packets to fill in the new trailing empty n rightmost slots in the feature vector. If there are not enough packets to fill the right most slots, we use zero padding to fill the remaining slots. For right shifts, we use the packet arrival time and packet size of the first index of the subsequence to fill in the vacated n slots.
@@ VI-A Feature Vector Construction
After training the encoder, the embedding generated by the encoder needs to be used for classification. Instead of directly using the embedding for clustering, we adopt an approach from an anomaly detection framework (AOC-IDS) [24]. In AOC-IDS, the average embedding of the normal traffic is computed, and the cosine similarity between the average and all of the embedding samples is fit to two Gaussian distributions (normal and abnormal traffic) using maximum likelihood estimation. During the binary classification, the test sample’s cosine similarity score is compared against the two distributions and labeled according to the distribution with a higher probability.
In our case, the encoder outputs similar embeddings for the same class, so we can characterize each sample based on how similar it is to the average embedding (centroid) of each class. Since our problem involves multiclass classification, we compute an average embedding for each of the ten classes in the training dataset. Then for a given sample, we compute its cosine similarity score to the 10 different average embeddings, which ultimately yields 10-dimensional cosine similarity score vector that is then used for clustering.
@@ VI-B GMM Clustering
To evaluate the impact of the 10-dimensional cosine similarity score vectors, we also run experiments where we cluster the 64-dimension embedding vectors directly from the same BiLSTM and FS-Net encoders. Table VI shows the result for the four different classification evaluations.
@@ VII Related Work
One of the state-of-the-art classifiers used in this paper is Flow Sequence Network (FS-Net) [7]. In FS-Net, packet size sequences are first passed through a small embedding layer, producing a learnable vector at each time step. The embeddings are then accepted by the GRU-based encoder and then subsequently fed into a GRU-based decoder. For classification, FS-Net combines the output of encoder and decoder and passes it through a dense layer to produce logits, which are used in the cross-entropy loss. The decoder’s output sequence is also compared against the original inputs to account for the reconstruction loss. The model is trained by minimizing the sum of the classification cross-entropy loss and the reconstruction cross-entropy loss.

### v4

@@ I Introduction
It is undesirable for background traffic to be labeled as an application traffic by a classifier, or for relevant application traffic to be incorrectly identified as background traffic. These misclassifications reduce both the performance and coverage needed for monitoring the usage, performance, and resource needs of different applications. In such scenarios, instead, it may be preferable that the deep learning model assign labels to traffic flow only when it can do so with high confidence. Therefore, in this paper, we develop a framework that can characterize the confidence of a model’s classification decision in a manner that significantly improves the classification performance. Our main innovations in this paper are as follows:
@@
    We investigate the trade-off between performance and coverage in traffic classification and subsequently design a classification framework with well-calibrated confidence scores that can help achieve different operating points in this trade-off based on the confidence threshold.
@@ III Deep Learning Classifiers: State of the Art
Various feature representations have been proposed for deep learning models. One of the earliest approaches simply fed the first $n$ raw bytes of each flow directly into a model for malware detection [21]. Building on this idea, another study [25] converted the first 512 bytes (beginning at the transport layer header) across 32 packets per flow into an image‐based representation for classification. As an alternative to feeding raw bytes of a flow directly into the model, ET-BERT [6] segmented traffic into “BURST,” which is a sequence of consecutive packets traveling in the same direction. Each BURST’s raw datagram bytes were split into two-byte bigrams, and the bigrams were converted to tokens for model input [6]. The model learned the contextual relationships among these tokens through self-supervised pre-training and was subsequently fine-tuned for different classification tasks [6]. However, one prior work has found that classifiers trained on raw packet payload bytes in encrypted traffic may overfit on unencrypted handshake fields. In their evaluation, masking SNI alone caused accuracy of the classifier to fall to 38.19%, and masking all handshake fields drove it down to 27.52% [10]. Supporting this concern, a different study [22] demonstrated that ET-BERT may overfit on specific packet-level information, ranging from SNI to session-specific header fields, when the dataset is not carefully curated. These findings highlight the importance of selecting robust features to ensure the model’s ability to generalize to traffic, where certain types of packet-level information may be unavailable.
In contrast, other studies employed time series features that are not dependent on specific packet bytes, such as sequences of packet sizes, arrival times, and directions for traffic classification [16, 7, 1]. Since these features are universally available in TCP and UDP traffic [^2] In this paper, we group UDP packets sharing the same 5-tuple and refer to them collectively as a UDP flow. and help prevent overfitting on packet level bytes [1], we experimentally evaluate classifiers that use time series features. We use Flow Sequence Network (FS-Net) [7] and our implementation of a Bidirectional LSTM (BiLSTM) [13], as FS-Net and various forms of LSTM have been used for evaluations in prior studies [1, 11, 23, 7, 6]. Each model is trained with cross-entropy loss, using a distinct combination of time series features and data balancing techniques. Although FS-Net overall achieves better performance, our BiLSTM yields comparable results, demonstrating that performance trends persist across different architectures, feature sets, and data balancing methods. Hence, we include both models in our evaluation. The architecture of our BiLSTM is as follows:
@@
We use the time series features defined in [1] to train the BiLSTM. Each flow is represented by a 40 × 2 feature vector corresponding to the first 40 packets, where at each index, the two features are: packet arrival time relative to the flow start and packet size. To encode directionality, we negate the packet size of the server packets. Since our training dataset has unequal number of flows for each class, we use data augmentation to balance the number of flows, since appropriate data augmentation improves training [20]. We adapt the translation technique that works well with time series features [20] and derive an augmented feature vector from a sampled original one as follows: we randomly select a start index of a subsequence and shift it either to the left or to the right by n steps. For left shifts, if the original traffic flow had more than 40 packets, we use the subsequent packets to fill in the new trailing empty n rightmost slots in the feature vector. If there are not enough packets to fill the right most slots, we use zero padding to fill the remaining slots. For right shifts, we use the packet arrival time and packet size of the first index of the subsequence to fill in the vacated n slots.
@@ VI-A Feature Vector Construction
After training the encoder, the embedding generated by the encoder needs to be used for classification. Instead of directly using the high dimensional (64-dim) embedding for clustering, we adopt an approach from an anomaly detection framework (AOC-IDS) that operates on cosine similarity score [24]. In AOC-IDS, the average embedding of the normal traffic is computed, and the cosine similarity between the average and all of the embedding samples is fit to two Gaussian distributions (normal and abnormal traffic) using maximum likelihood estimation. During the binary classification, the test sample’s cosine similarity score is compared against the two distributions and labeled according to the distribution with a higher probability.
In our case, the encoder outputs similar embeddings for the same class, so we can characterize each sample based on how similar it is to the average embedding (centroid) of each class. Since our problem involves multiclass classification, we compute an average embedding for each of the ten classes in the training dataset. Then for a given sample, we compute its cosine similarity score to the 10 different average embeddings, which yields a compact 10-dimensional cosine similarity score vector.
@@ VI-B GMM Clustering
To evaluate the impact of the 10-dimensional cosine similarity score vectors, we also run experiments where we cluster the 64-dimensional embedding vectors directly from the same BiLSTM and FS-Net encoders. Table VI shows the result for the four different classification evaluations.
@@ VII Related Work
One of the reference model used in this paper is Flow Sequence Network (FS-Net) [7]. In FS-Net, packet size sequences are first passed through a small embedding layer, producing a learnable vector at each time step. The embeddings are then accepted by the GRU-based encoder and then subsequently fed into a GRU-based decoder. For classification, FS-Net combines the output of encoder and decoder and passes it through a dense layer to produce logits, which are used in the cross-entropy loss. The decoder’s output sequence is also compared against the original inputs to account for the reconstruction loss. The model is trained by minimizing the sum of the classification cross-entropy loss and the reconstruction cross-entropy loss.